the ratio of an angle to its complement is 4:5. find the measure of each angle?
Answer by palanisamy(496)

Let the required angle be x.
The its complementary angle = 90-x
Given, x/(90-x) = 4/5
5x = 4*(90-x)
5x = 360-4x
5x+4x = 360
9x = 360
x = 360/9
x = 40degrees
Its complementary angle = 90-40 = 50 degrees